OpenStudy (anonymous):

the volume of a cylinder is 328cm3. find the radius of the base if the cylinder has a height of 8cm. round your answer to the nearest 0.1cm can someone help me

jhonyy9 (jhonyy9):

are of the base is pir*2 =3,14 x r*2 the vol. is are base x height = 3,14xr*2x8=328 -- -- 25,12xr*2=328 --- r*2=328/25,12 --- r=(328/25,12)*(1/2)